OpenStudy (anonymous):

integrate (1-cos x) / (1+cos x)

hartnn (hartnn):

do u know the half angle formulas ? cos 2x = ?

OpenStudy (anonymous):

I got it

hartnn (hartnn):

u know these ? cos 2x= 1-2 sin^2 x = 2cos^2x-1 ?

hartnn (hartnn):

oh,ok. :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1-cosx=2sin^2(x/2) 1+cosx=2cos^2(x/2)

OpenStudy (anonymous):

thanks for the help